# Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale

### v2

@@ Abstract
Social engineering attacks delivered via email, commonly known as phishing, represent a persistent cybersecurity threat leading to significant organizational incidents and data breaches. Although many organizations train employees on phishing, often mandated by compliance requirements, the real-world effectiveness of this training remains debated.
@@
To contribute to evidence-based cybersecurity policy, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm. Our experimental design refined prior work by comparing training modalities in operational environments, validating NIST’s standardized phishing difficulty measurement, and introducing novel organizational-level temporal resilience metrics. Echoing prior work, training interventions showed no significant main effects on click rates (pp=0.450) or reporting rates (pp=0.417), with negligible effect sizes. However, we found that the NIST Phish Scale predicted user behavior, with click rates increasing from 7.0% for easy lures to 15.0% for hard lures. Our organizational-level resilience result was mixed: 36-55% of campaigns achieved “innoculation” patterns where reports preceded clicks, but training did not significantly improve organizational-level temporal protection. In summary, our results confirm the ineffectiveness of current phishing training approaches while offering a refined study design for future work.
1 Introduction
@@
Phishing attacks harm organizations through operational incidents, data breaches, and financial losses [4]. Industry studies and the US FBI agree that phishing is the primary attack vector to breach organizations [42]. As an example, in 2022–2023 campaigns attributed to the Scattered Spider cluster (aka 0ktapus/Muddled Libra) compromised 9,931 accounts in over 130 organizations [43, 23]. This campaign cost MGM Resorts alone an estimated \$100 million [79]. To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training [58, 48].
@@
Despite widespread organizational investment in anti-phishing training, mounting empirical evidence (and employee sentiment [78]) questions its effectiveness. Large-scale studies across multiple sectors demonstrate minimal training impact [48, 53, 54]. Most pertinently, Lain et al. showed that phishing training did not improve security outcomes across 19,000+ employees over multiple years [53, 54]. These findings align with theoretical work by Volkamer et al., who demonstrate fundamental limitations in human processing of security indicators [96].
@@
This growing body of evidence suggests that phishing training is ineffective, but we believe that further reproduction of the result is needed to inform cybersecurity policy and defense strategies [67, 89]. [^1] Our reproduction effort differs from an exact replication, which would involve repeating identical procedures. Instead we focus on testing the generalizability of the phenomenon of training ineffectiveness, varying the organizational context and the training approach. We target two specific knowledge gaps. First, existing studies lack standardized measures of phishing difficulty, so we do not know whether training effectiveness varies by attack sophistication. Second, most studies examine single training modalities rather than comparing different approaches systematically, so we do not know if some training designs are more effective than others.
@@
To address these limitations, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm. Our study design has three distinctive elements: (1) mandatory pre-phishing training using two distinct training modalities (lecture-based vs. interactive); (2) systematic difficulty assessment using the NIST Phish Scale [86, 7]; and (3) novel organizational temporal resilience measurement through the Organizational Inoculation Index, which measures the effectiveness of phishing training from an organizational perspective rather than an individual one. This design enables rigorous evaluation of training effectiveness across standardized difficulty levels while providing new insights into organizational cybersecurity resilience patterns that complement individual-focused assessments.
@@
Our reproduction study confirms the ineffectiveness of anti-phishing training across multiple dimensions, but adds nuance to prior results. Consistent with prior work, neither training modality produced statistically significant improvements in phish click reduction (pp=0.450) or phish reporting behavior (pp=0.417), with negligible effect sizes indicating minimal practical value for organizational security investment. However, the NIST Phish Scale successfully predicted user behavior (F(2, 12086) = 41.415, p < 0.001, $\eta^{2}$ = 0.007), providing the first large-scale validation of this standardized difficulty measure. [^2] While the effect size is small in statistical terms due to low base rates of phishing engagement, the practical implications are significant: click rates doubled from 7.0% to 15.0% as lure difficulty increased. Beyond that, our Organizational Inoculation Index revealed substantial organizational resilience that operates independently of individual training effectiveness. Across all experimental conditions, 36.4% to 54.5% of template-group combinations demonstrated “inoculation” patterns where threat reports (from a security-conscious employee) preceded threat interaction by another employee.
We summarize our main contributions below:
  1. 1.
    Reproduction of Training Ineffectiveness Across Multiple Modalities. Our results reproduce the core finding of training ineffectiveness while demonstrating this phenomenon’s generalizability across training designs.
  2. 2.
@@
    First Large-Scale Validation of the NIST Phish Scale. We evaluate the NIST Phish Scale across 12,511 participants. The Phish Scale is a viable measure for assessing phishing difficulty, with predictive utility.
  3. 3.
@@
    Introduction of the Organizational Inoculation Index. We introduce and validate a novel metric for organizational temporal security. We report substantial organizational resilience (36-55% inoculation rates), which operates independent of individual training effectiveness and provides new insights into collective cybersecurity behaviors.
@@
Significance: Cybersecurity practice and policy must be evidence-based. For regulators, our work contributes to growing evidence that compliance-driven training requirements may not achieve intended security outcomes. For organizations, our results suggest realistic (low) expectations about training outcomes and highlight the importance of technical controls rather than human-centered defenses. However, we also demonstrated that collective security behaviors create measurable protective effects, emphasizing the value of feedback loops between user reports and real-time threat response. For researchers, we validate NIST’s standardized difficulty measure and confirm the reproducibility of training ineffectiveness findings across organizational contexts.
2 Background and Related Work
Figure 1: Phishing defense architecture, mapping the email pathway from authentication through post-delivery response. Technical controls create multiple interception points prior to human interaction, and human reports provide a feedback opportunity for organizational-level threat mitigation. See Appendix 2 for the full control pathway in the studied organization.
@@ 2 Background and Related Work
We review phishing attacks and defenses (§2.1), and then discuss prior empirical studies on phishing training (§2.2).
2.1 Phishing Attacks and Defenses
@@ 2.1.1 Phishing Attacks
In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests, and the interests of their organizations. Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine [4, 34]. Traditionally, phishing attacks are divided into two classes [25]:
Mass Phishing Campaigns: These use malicious emails sent to thousands or millions of recipients, with common themes such as banking, shipping, and software updates [4]. Individual success rates are low (1-3%) [3], but the massive scale makes them profitable.
Spear Phishing and Targeted Attacks: Spear phishing targets specific individuals or organizations with personalized content from open-source intelligence (OSINT) and reconnaissance. Attackers mine sources such as social media to craft messages that reference real relationships and events [13]. These tailored campaigns have higher success rates (10–20%) [14]. Successful spear phishing attacks include the 2016 Democratic National Committee breach and widespread business email compromise (BEC) schemes [4].
Modern Phishing: Contemporary phishing operations leverage infrastructure including compromised (“trusted”) websites and domain generation algorithms to maintain persistence and evade takedown efforts [3]. In addition, Generative AI is enabling both mass phishing and spear phishing through targeted and fluent emails and other media [4, 80]. LLM-generated phishing content now achieves effectiveness comparable to human experts in organizational settings [9].
2.1.2 Automated Defenses
Figure˜1 illustrates how these attacks flow through an organization’s email system. Many technical defenses (blue boxes) combat attackers’ behavior, such as SPF, DKIM, and DMARC to verify that mail comes unaltered from authorized infrastructure [51, 22, 52]; ARC for preserving authentication results across forwarding [5]; BIMI for verified logos [10, 12]; and secure email gateways (SEG) for content and header inspection and URL rewriting for time-of-click inspection [1].
After delivery, organizations employ continuous monitoring and feedback loops to adapt controls to evolving campaigns. Organizations blend threat-intelligence updates with AI/ML-driven anomaly scoring and impersonation detection to surface novel lures [58]. For example, security operations center (SOC) analysts triage user-reported messages, promote high-confidence indicators back into blocklists, and tune SEG policies based on observed false positives/negatives. Because adversaries adapt, detection is a dynamic process [4, 3].
See Appendix 2 for the details of the studied organization’s email controls and how user phishing reports create feedback loops that strengthen automated defenses.
2.1.3 Phishing Training as a Defense
@@
Because phishing attacks compromise these technical defenses [14, 13], the human recipients play an important role in phishing defense [30, 15, 2, 28]. There are many approaches to training humans about phishing attacks. Methods range from lecture-style training involving passive videos and readings [56] to simulation feedback providing immediate feedback after interaction with phishing simulations [19, 18]. Interactive training uses gamification or scenario-based learning to engage participants [11, 46], while just-in-time training alerts users in real time during risky behavior [36].
@@
Anti-phishing trainings are often included in the cybersecurity awareness trainings ubiquitous in organizations. Legal mandates drive this practice, e.g., HIPAA requiring security awareness training for healthcare organizations [74]. GDPR and PCI DSS encourage or mandate similar training [76, 21]. Industry standards and frameworks further reinforce these requirements [75], e.g., ISO 27001 calling for human resource security controls including awareness programs [35].
2.2 Empirical Studies of Phishing Training
Table 1: Summary of Prior Experiments on Phishing and Training Interventions. Studies are organized by research context (Lab vs. Real-World). Context: “Lab” denotes research simulations, where subjects are not in their work roles. “Real-World” denotes studies where participants were employees acting in their job capacity. Complexity measure: Whether the study controlled for the difficulty of the phishing tasks. indicates the use of an open control such as NIST Phish Scale, and indicates no control. Hypotheses: How these works informed the hypotheses in the present study (cf. Table˜4).
[82] (2024) Lab University students 96 H3
[16] (2024) Lab University students 117 (Phish Scale) H4
[18] (2014) Real-world Employees 1,359 H2, H4
[19] (2017) Real-world University students 150 H2, H3
[90] (2022) Real-world Employees 119 H2
[53] (2022) Real-world Employees 19,000+ H2
[47] (2023) Real-world Finance employees 5,000 H2
[54] (2024) Real-world Employees 4,554 H2, H3
[29] (2024) Real-world Banking employees “Thousands” H2
[48] (2025) Real-world Healthcare employees 19,500+ H2
Ours Real-world Fintech org. 12,511 (Phish Scale)
@@ 2.2 Empirical Studies of Phishing Training
Our review and analysis of prior work is summarized in Table˜1. We contribute to the body of knowledge on phishing training by conducting another large-scale study situated in a real-world context (table columns 2-4), incorporating a complexity measure (column 5), and a comparison of multiple training approaches (column 6). Our study tests hypotheses derived from prior work (final column). The remainder of this section describes these contributions. For general surveys of knowledge about human factors in phishing training, we refer the reader to Marshall et al. [56] and Franz et al. [37].
@@
Lab vs. Real-world Studies: Many early phishing training studies were conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control. Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes [90, 56]. However, as noted by Franz et al. [37], these controlled settings may lack ecological validity. In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students.
@@
In response, a growing number of studies have begun examining phishing training in real-world organizational environments. For example, Doing et al. [29] analyzed phishing behaviors in bank employees, and Ho et al. [48] studied healthcare employees. Lain et al. has also conducted large-scale real-world measurements of training effectiveness [53, 54] Our study echoes some of these prior results in a different organizational context.
@@
Controlling for Lure Complexity: These studies consistently report little effectiveness of anti-phishing training. These works have not, however, incorporated a standardized method for assessing the difficulty of the phishing lures they use [48, 29]. We contribute the first real-world measurement of the effect of lure complexity on subjects’ interaction behaviors, based in NIST’s Phish Scale.
@@
To assess phishing lure complexity, the NIST Phish Scale categorizes them along two dimensions (Table˜2). Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [16]. Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [6]. Higher alignment messages with fewer cues are the most deceptive [14]. The premise of Phish Scale is to allow organizations to benchmark lure complexity and contextualize simulation outcomes [17]. Barrientos et al. [8] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [27] provided the official implementation guidelines. Ours is the first study to validate the Phish Scale’s premise at scale.
@@
Training Approaches: Franz et al. [37] taxonomize five categories of phishing interventions: training, awareness, education, notifications, and design interventions. Most studies have focused on training, which can be divided into lecture-style annual training and embedded training (e.g., through ongoing phishing simulations sent by the organization’s security team). Like some prior studies, our study performs a controlled experiment to compare the effectiveness of two forms of phishing training.
Table 2: NIST Phish Scale for Detection Difficulty [87, 27].
Cues \ Alignment Weak Medium Strong
Few Medium Hard Hard
Some Easy-to-Medium Medium Hard
Many Easy Medium Medium
@@
This study addresses those gaps by applying the NIST Phish Scale [27] to evaluate training effectiveness across varied phishing difficulty levels in an enterprise deployment with over 12,000 participants. In this study, our objectives are to: (1) operationalize the NIST framework at scale, (2) examine how difficulty influences detection and reporting across training modalities, (3) provide comparative evidence on lecture-based versus interactive approaches, and (4) offer guidance for large-scale awareness efforts under conditions of modest effectiveness [48]. This enables evidence-based decision-making on phishing defense strategies that account for both training design and message complexity.
3 Methodology
To summarize the preceding analysis: phishing training is widely practiced, though recent evidence sheds doubt on its utility. Few studies have rigorously assessed outcomes against standardized difficulty measures at sufficient scale for robust inference. We provide a reproduction of this observation, tested through hypotheses derived from prior works.
In this section, we describe our methodology: study design, subjects, hypotheses, metrics, and analysis methods. For ethics considerations, see the associated appendix.
3.1 Study Design

### v3

# Anti-Phishing Training (Still) Does Not Work: A Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale

@@ Abstract.
Social engineering attacks delivered via email, commonly known as phishing, represent a persistent cybersecurity threat leading to significant organizational incidents and data breaches. Although many organizations train employees on phishing, often mandated by compliance requirements, the real-world effectiveness of this training remains debated. Past work has demonstrated the ineffectiveness of training, but reproduction across different organizations, training approaches, and with a standardized threat assessment will help the generalizability of this phenomenon.
@@
To contribute to evidence-based cybersecurity policy, we conducted a large-scale reproduction study (NN=12,511) at a US-based financial technology firm. Our design refined prior work by comparing training modalities in operational environments, applying NIST’s standardized phishing difficulty measurement, and introducing novel organizational-level temporal resilience metrics.
Echoing prior work, training interventions showed no significant main effects on click rates (pp=0.450) nor reporting rates (pp=0.417), with negligible effect sizes. However, we found that the NIST Phish Scale predicted user behavior, with click rates increasing from 7.0% (easy lures) to 15.0% (hard lures). Our organizational-level resilience result was mixed: 36-55% of campaigns achieved “inoculation” patterns where reports preceded clicks, but training did not significantly improve organizational-level temporal protection. Our results confirm the ineffectiveness of current phishing training approaches and offer a refined study design for future work.
copyright: none
1. Introduction
@@
Phishing is a primary attack vector (greitzerExperimentalInvestigationTechnical2021, ) and regularly harms organizations through operational incidents, data breaches, supply chain attacks(valsorda-compromise-survey, ),and financial losses (alkhalilPhishingAttacksRecent2021, ). For, in 2022–2023 campaigns attributed to the Scattered Spider cluster compromised 9,931 accounts in over 130 organizations (groupib2022oktapus, ; cisa2023scatteredspider, ). This campaign cost MGM Resorts alone an estimated \$100 million (reuters2024mgm100m, ). To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training (mengesContrastingSynergizingCISOs2024, ; hoUnderstandingEfficacyPhishing2025, ).
@@
Despite organizational and regulatory emphasis on anti-phishing training, its effectiveness is questioned by mounting empirical evidence (hoUnderstandingEfficacyPhishing2025, ; lain2022phishing, ; lain2024content, ) and employee sentiment (reevesGetRedhotPoker2021, ). Most pertinently, Lain et al. found that phishing training did not improve security outcomes across 19,000+ employees over years (lain2022phishing, ; lain2024content, ). These findings align with theoretical work by Volkamer et al., who show fundamental limitations in human processing of security indicators (volkamer2020making, ).
@@
These prior studies support the general contention that phishing training is ineffective. In this work we refine our understanding by addressing two specific knowledge gaps. First, existing studies lack standardized measures of phishing difficulty, so we do not know whether training effectiveness varies by attack sophistication. Second, most studies examine single training modalities rather than comparing different approaches systematically, so we do not know if some training designs are more effective than others.
@@
To address these limitations, we conducted a large-scale reproduction study (NN=12,511) at a US-based financial technology firm. [^1] Our reproduction effort differs from replication, which would repeat identical procedures. Instead, we focus on testing the generalizability of the phenomenon of training ineffectiveness, varying organizational context and training approach. Our study design has three distinctive elements: (1) mandatory pre-phishing training using two distinct training modalities (lecture-based vs. interactive); (2) systematic difficulty assessment using the NIST Phish Scale (stevesCategorizingHumanPhishing2020, ; barrientosScalingPhishAdvancing2021, ); and (3) a novel organizational temporal resilience measurement through our Organizational Inoculation Index. Our design evaluates training effectiveness across standardized difficulty levels, and brings new insights into organizational cybersecurity resilience patterns to complement individual ones.
@@
We confirm the ineffectiveness of anti-phishing training across multiple dimensions, but adds nuance to prior results. Consistent with prior work, neither training modality produced statistically significant improvements in phish click reduction (pp=0.450) nor phish reporting behavior (pp=0.417), with negligible effect sizes indicating minimal practical value for individuals. However, the NIST Phish Scale successfully predicted user behavior (F(2, 12086)=41.415, p ¡ 0.001, $\eta^{2}$=0.007), providing the first large-scale validation of this standardized difficulty measure. [^2] The effect size is (statistically) small due to low base rates of phishing engagement. Practically, click rates doubled from 7.0% to 15.0% as lure difficulty increased. Beyond that, our Organizational Inoculation Index revealed substantial organizational resilience that operates independently of individual training effectiveness. Across all experimental conditions, 36%–55% of template-group combinations demonstrated “inoculation” patterns where threat reports from one employee preceded threat interaction by others. In summary, we contribute:
  1. (1)
    Reproduction of Training Ineffectiveness. We reproduce the core finding of phishing training ineffectiveness.
  2. (2)
@@
    First Large-Scale Validation of NIST Phish Scale. We show the NIST Phish Scale predicts phishing lure difficulty.
  3. (3)
@@
    Introduction of the Organizational Inoculation Index. We introduce a novel metric for organizational temporal security. We find substantial organizational resilience (36-55% inoculation rates), independent of individual training effectiveness.
@@
Significance: Cybersecurity practice and policy must be evidence-based. For regulators, our work adds more evidence that phishing training is ineffective. For organizations, our results suggest realistic (low) expectations about training outcomes and highlight the importance of technical controls to assist humans. We also found that collective security behaviors may create measurable protective effects, emphasizing the value of user/IT feedback loops for real-time threat response. For researchers, we validate NIST’s standardized difficulty measure and confirm the reproducibility of training ineffectiveness findings across organizational contexts.
2. Background and Related Work
Table 1. Summary of Prior Experiments on Phishing and Training Interventions. Studies are organized by research context (Lab vs. Real-World). Context: “Lab” denotes research simulations, where subjects are not in their work roles. “Real-World” denotes studies where participants were employees acting in their job capacity. Complexity measure: Whether the study controlled for the difficulty of the phishing tasks. indicates the use of an open control such as NIST Phish Scale, and indicates no control. Hypotheses: How these works informed the hypotheses in the present study (cf. Table 3).
(schoniYouKnowWhat2024, ) (2024) Lab University students 96 H3
(canhamNotAllVictims2024, ) (2024) Lab University students 117 (Phish Scale) H4
(caputoGoingSpearPhishing2014, ) (2014) Real-world Employees 1,359 H2, H4
(carellaImpactSecurityAwareness2017, ) (2017) Real-world University students 150 H2, H3
(sumnerExaminingFactorsImpacting2022, ) (2022) Real-world Employees 119 H2
(lain2022phishing, ) (2022) Real-world Employees 19,000+ H2
(hillmanEvaluatingOrganizationalPhishing2023, ) (2023) Real-world Finance employees 5,000 H2
(lain2024content, ) (2024) Real-world Employees 4,554 H2, H3
(doingAnalysisPhishingReporting2024, ) (2024) Real-world Banking employees “Thousands” H2
(hoUnderstandingEfficacyPhishing2025, ) (2025) Real-world Healthcare employees 19,500+ H2
Ours Real-world Fintech org. 12,511 (Phish Scale)
@@ 2. Background and Related Work
We review phishing attacks and defenses (§2.1), and prior empirical studies on phishing training (§2.2).
2.1. Phishing Attacks and Defenses
@@ 2.1.1. Phishing Attacks
In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests, and the interests of their organizations. Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine (alkhalilPhishingAttacksRecent2021, ; ferreiraPersuasionHowPhishing2019, ). These attacks use compromised infrastructure, browser vulnerabilities(10.1145/3589334.3645535, ), look-alike domains, legitimate certificates (kim2021sooel, ) and social engineering to evade technical defenses (10.1145/3696410.3714678, ; 10.1145/3652963.3655042, ) and exploit human trust (alkhalilPhishingAttacksRecent2021, ; aleroudPhishingEnvironmentsTechniques2017, ). Traditionally, phishing attacks are divided into two classes (dasSoKComprehensiveReexamination2020, ): mass phishing (form emails sent to many targets, often with a phishing kit (10.1145/3696410.3714710, )) and spear phishing (custom emails sent to specific targets).
2.1.2. Automated Defenses
Many technical defenses combat attackers’ behavior, such as SPF, DKIM, and DMARC to verify that mail comes unaltered from authorized infrastructure (rfc7208spf2014, ; rfc6376dkim2011, ; rfc7489dmarc2015, ); ARC for preserving authentication results across forwarding (rfc8617arc2019, ); BIMI for verified logos (bimi-group, ; bimi-ietf-guidance-2025, ); and secure email gateways (SEG) for content and header inspection and URL rewriting for time-of-click inspection (msft-safe-links-2025, ). After delivery, organizations employ continuous monitoring and feedback loops to adapt controls to evolving campaigns (mengesContrastingSynergizingCISOs2024, ). For an example of the dynamics (alkhalilPhishingAttacksRecent2021, ; aleroudPhishingEnvironmentsTechniques2017, ), security operations center (SOC) analysts triage user-reported messages, promote high-confidence indicators back into blocklists, and tune SEG policies based on observed false positives/negatives.
Appendix A describes elements of the studied organization’s email controls and feedback loops.
2.1.3. Phishing Training as a Defense
@@
Because phishing attacks compromise these technical defenses (butaviciusBreachingHumanFirewall2015, ; burnsSpearPhishingBarrel2019, ), the human recipients play an important role in phishing defense (downsDecision2006, ; canfieldBetter2019, ; abroshanPhishingHappensTechnology2021, ; desoldaHuman2022, ). Training methods range from lecture-style training involving passive videos and readings (marshallExploringEvidenceEmail2024, ) to simulation feedback providing immediate feedback after interaction with phishing simulations (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ). Interactive training uses gamification or scenario-based learning to engage participants (bitrianGamificationWorkforceTraining2024, ; hartRiskioSeriousGame2020, ), while just-in-time training alerts users in real time during risky behavior (franklinOptimisingNudgesBoosts2019, ).
@@
Anti-phishing trainings are often included in the cybersecurity awareness trainings ubiquitous in organizations. Legal mandates drive this practice. For example: HIPAA requires security awareness training for healthcare organizations (hipaa-security-rule, ), and GDPR and PCI DSS encourage or mandate similar training (gdpr-article39, ; pci-dss, ). Industry standards reinforce these requirements (nist-csf, ), e.g., ISO 27001 calling for human security controls like awareness programs (iso-27001, ).
2.2. Empirical Studies of Phishing Training
@@
Our review and analysis of prior work is summarized in Table 1. We contribute to the body of knowledge on phishing training by conducting another large-scale study situated in a real-world context (table columns 2-4), incorporating a complexity measure (column 5), and a comparison of multiple training approaches (column 6). Our study tests hypotheses derived from prior work (final column). The remainder of this section describes these contributions. For general surveys of knowledge about human factors in phishing training, we refer the reader to (marshallExploringEvidenceEmail2024, ; franzSoK, ).
@@
Lab vs. Real-world Studies: Many early phishing training studies were conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control. Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes (sumnerExaminingFactorsImpacting2022, ; marshallExploringEvidenceEmail2024, ). However, as noted by Franz et al. (franzSoK, ), these controlled settings may lack ecological validity. In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students.
@@
In response, a growing number of studies have begun examining phishing training in real-world organizational environments. For example, Doing et al. (doingAnalysisPhishingReporting2024, ) analyzed phishing behaviors in bank employees, and Ho et al. (hoUnderstandingEfficacyPhishing2025, ) studied healthcare employees. Lain et al. has also conducted large-scale real-world measurements of training effectiveness (lain2022phishing, ; lain2024content, ) Our study echoes some of these prior results in a different organizational context.
@@
Controlling for Lure Complexity: These studies consistently report little effectiveness of anti-phishing training. These works have not, however, incorporated a standardized method for assessing the difficulty of the phishing lures they use (hoUnderstandingEfficacyPhishing2025, ; doingAnalysisPhishingReporting2024, ).
@@
The US NIST has proposed a scale for this purpose (stevesCategorizingHumanPhishing2020, ; dawkinsScalingNISTPhish2023, ). We summarize it here and give more details in Appendix B. The NIST Phish Scale categorizes lure difficulty along two dimensions. Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities (canhamNotAllVictims2024, ). Premise Alignment measures the relevance of the email to the recipient’s organizational context (antoniouSystematicMethodExecute2015, ). Higher alignment messages with fewer cues are the most deceptive (butaviciusBreachingHumanFirewall2015, ). The premise of Phish Scale is to allow organizations to benchmark lure complexity and contextualize simulation outcomes (canhamPhishDerbyShoring2022, ). Building on early validation work (barrientosScalingPhishAdvancing2021, ; dawkinsScalingNISTPhish2023, ), we contribute the first real-world measurement of the effect of lure complexity on subjects’ interaction behaviors.
@@
Training Approaches: Franz et al. (franzSoK, ) taxonomize five categories of phishing interventions: training, awareness, education, notifications, and design interventions. Most studies have focused on training, which can be divided into lecture-style annual training and embedded training (e.g., via ongoing phishing simulations from the organization’s security team). Like a few prior studies (caputoGoingSpearPhishing2014, ; carellaImpactSecurityAwareness2017, ; lain2024content, ), we compare the effectiveness of two forms of phishing training.
@@
This study addresses those gaps by applying the NIST Phish Scale (dawkinsScalingNISTPhish2023, ) to evaluate training effectiveness across varied phishing difficulty levels in an enterprise deployment with over 12,000 participants. In this study, our objectives are to: (1) operationalize the NIST framework at scale, (2) examine how difficulty influences detection and reporting across training modalities, (3) provide comparative evidence on lecture-based versus interactive approaches, and (4) offer guidance for large-scale awareness efforts under conditions of modest effectiveness (hoUnderstandingEfficacyPhishing2025, ). This enables evidence-based decision-making on phishing defense strategies that account for both training design and message complexity.
3. Methodology